# CricaVPR: Cross-image Correlation-aware Representation Learning for Visual Place Recognition

### v1

@@ 4.2 Implementation Details
We fine-tune our model on two NVIDIA GeForce RTX 3090 GPUs using PyTorch. The resolution of the input image is 224×224 and the token dimension of the backbone (ViT-B/14) is 768. Our model outputs the 14×768-dim original global features, and we use PCA for dimensionality reduction. The bottleneck ratio of our MulConv adapters is set to 0.5, so the input dimension of the three convolutional paths is 384. The 1×1 convolution before the 3×3 and 5×5 convolution reduces the channel to 24. The output dimensions of the three convolutional paths are 192, 96, and 96. We set the hyperparameters $\alpha=1,\beta=50,\lambda=0$ in Eq. 5 and margin = 0.1 in online mining, as in GSV-Cities [1]. We fine-tune the model using the Adam optimizer with the initial learning rate set as 0.0001 and multiplied by 0.5 after every 3 epochs. A training batch contains 72 places with 4 images each (i.e. 288 images). Training is implemented until the R@5 on Pitts30k does not improve for 3 epochs. An inference batch contains 8 images for Pitts30k and 16 images for others.
4.3 Comparisons with State-of-the-Art Methods
In this section, we compare our CricaVPR with several SOTA VPR methods, mainly including six global-retrieval-based methods: NetVLAD [5], SFRS [23], CosPlace [8], GCL [38], MixVPR [2] and EigenPlaces [10]. Note that our work uses the same training dataset as MixVPR, i.e., GSV-Cities. Meanwhile, CosPlace and EigenPlaces are trained on individually constructed extra large-scale datasets, i.e., SF-XL. Both MixVPR and EigenPlaces are the latest works and represent the SOTA performance of the VPR methods based on global feature retrieval. Additionally, we also compare our approach with two excellent two-stage VPR methods (Patch-NetVLAD [26] and TransVPR [61]), which require time-consuming re-ranking using local features. The details of these methods are in Suppl. Material. Table 2 shows the quantitative results on Pitts30k, Tokyo24/7, and MSLS. Our CricaVPR method achieves the best R@1/R@5/R@10 on all datasets.
@@ 4.3 Comparisons with State-of-the-Art Methods
To further evaluate the performance of our method in extreme scenarios, we conduct experiments on three challenging datasets: Nordland, which exhibits seasonal changes; AmsterTime, which spans a very long time period; and SVOX, which shows extreme illumination and weather variations. The results, as shown in Table 3, demonstrate the significant superiority of our method compared to other SOTA methods. Our CricaVPR outperforms all other SOTA methods with 14.5%, 15.8%, and 20.7% absolute R@1 improvements on Nordland, AmsterTime, and SVOX-Night, respectively. This further highlights that our global representation is highly robust.
Moreover, our model only costs 3.5 hours for training, which is significantly less than the full-day time used by CosPlace/EigenPlaces. The training epochs of our model (10 epochs) are also less than MixVPR (30 epochs) using the same dataset. More details are in Suppl. Material.
@@ 4.4 Ablation Study
Table 4: Ablation on cross-image awareness. The “+Crica” represents the addition of our cross-image correlation awareness to get the final global feature. The “SPM” represents our spatial pyramid model representation, while “SPMG” is the spatial pyramid model solely based on GeM. Except for the DINOv2-GeM and DINOv2-SPM that directly use an untuned backbone (as baseline), all other versions use our adaptation method for fine-tuning.
@@
Ablation on cross-image correlation awareness. The cross-image correlation awareness achieved by the cross-image encoder after backbone is the most important module in our method. We compare the performance of the three kinds of global features with or without across-image awareness. These features are GeM, the spatial pyramid model representation solely based on GeM (SPMG), and our spatial pyramid model representation (SPM) which uses both class token and GeM. The results are shown in Table 4. After incorporating the proposed cross-image correlation awareness (Crica), all three features achieve significant performance improvements. Due to the already impressive performance of the SPM feature after our model adaptation (AdaptSPM), the improvement provided by Crica on this feature is not as pronounced as on the GeM and SPMG features. Nevertheless, AdaptSPM+Crica still resulted in 4.2%, 7.9%, and 4.4% absolute R@1 improvements over the AdaptSPM feature on Pitts30k, Tokyo24/7, and MSLS-val, respectively. Moreover, AdaptGeM+Crica achieves an impressive 17.4% absolute R@1 improvement over AdaptGeM on Tokyo24/7. With the combined effect of our Crica and model adaptation, our method achieves nearly 2× higher R@1 on Tokyo24/7 and MSLS-val compared to the direct use of frozen DINOv2 with the SPM representation (FrozenDINOv2-SPM). In the next subsection, we will further validate our model adaptation.
Ablation on adaptation. We first use only the GeM features alone (without cross-image awareness) to demonstrate the performance improvement achieved by our adaptation method. As shown in Table 5, MulConvAdapter-GeM using our adaptation achieves a significant improvement over FrozenDINOv2-GeM. Especially on MSLS-val, which has more dynamic interference, our adaptation achieves nearly 2× higher R@1. Fig. 6 vividly illustrates the underlying reasons. The adapted DINOv2, in contrast to the pre-trained DINOv2, exhibits a greater ability to attend to objects that are relevant for place recognition, with more attention given to more important objects. Table 5 also shows the performance of different fine-tuning methods when using the proposed global features (i.e. the SPM feature with cross-image correlation awareness). The FullTunedDINOv2 achieves a notable improvement over FrozenDINOv2 on Pitts30k and MSLS-val. However, because our training data has no night images like those in Tokyo24/7, FullTunedDINOv2 performs worse than FrozenDINOv2 on Tokyo24/7, i.e., it suffers from catastrophic forgetting. This indicates the necessity of parameter-efficient fine-tuning (using an adapter). Besides, ConvAdapter (as in Convpass [31]) uses 3×3 convolution to introduce local inductive biases into the model. However, it brings inappropriate local priors for VPR and results in performance degradation compared to VanillaAdapter. Our method (MulConvAdapter) uses multi-scale convolution to introduce more proper local information and thus achieves the best performance.
@@
Figure 6: The output feature map (attention) visualizations of pre-trained DINOv2 and adapted DINOv2. The regions attended to by pre-trained DINOv2 have no relevance to place recognition. However, adapted DINOv2 focuses on discriminative areas for VPR. Buildings that remain relatively unchanged over time receive the highest attention. Vegetation that is not expected to change in the short term receives moderate attention. Non-discriminative elements such as the sky and ground, as well as dynamic vehicles, are ignored.
@@
Ablation on dimensions of our descriptor. Intuitively, higher-dimensional descriptors have the ability to provide more information, thus offering the potential to achieve better performance. However, removing redundant or noisy information often leads to performance improvements as well. We analyze the impact of descriptor dimensionality on performance, and the results are shown in Table 6. Our method gets the best performance when using PCA to reduce the descriptor dimension to 4096-dim, so it is the default dimensionality we recommend. Furthermore, we continue to reduce the dimensionality to observe the point at which performance starts to noticeably decline on each dataset. For Pitts30k, the 512-dim descriptor still achieves an impressive 94.5% R@1, with no significant decrease compared to the 4096-dim descriptor. However, using the 512-dim descriptor on the other two datasets results in an obvious performance drop. This is mainly due to the drastic condition changes and perceptual aliasing issue in these datasets, requiring higher-dimensional descriptors to provide sufficient information to distinguish places. When there is a pressing need for low-dimensional descriptors, we suggest using the 512-dim descriptor for Pitts30k, while selecting the 1024-dim or 2048-dim descriptor for Tokyo24/7 and MSLS.
@@
5 Conclusions
In this paper, we presented CricaVPR, a robust global representation method with cross-image correlation awareness for VPR. Our method leverages the cross-image encoder equipped with self-attention to establish the correlation among multiple images within a batch, enabling the model to harvest useful information from other images while generating the feature representation of an image. This makes the produced global features condition-invariant, viewpoint-invariant, and capable of addressing perceptual aliasing. Furthermore, we proposed a multi-scale convolution-enhanced adaptation method to introduce proper local information and effectively unleash the capability of the pre-trained foundation model for VPR. Experimental results on several VPR benchmark datasets demonstrate that our CricaVPR can provide a robust global representation to address various challenges in VPR and outperforms SOTA methods by a significant margin.
Acknowledgments
Supplementary Material
6 Overview
7. Training Epoch, Training Time & Tunable Parameters
8. Visualizations of Place Features Using t-SNE
9. Additional Results on Challenging Datasets
10. Additional Ablations on Cross-image Encoder
11. Effects of Training Batch Size
12. Additional Ablations on MulConvAdapter
13. Effects of Adaptation on the Used SPM Feature
14. Datasets Details
15. Compared Methods Details
16. Additional Qualitative Results and Failure Cases
7 Training Epoch, Training Time, and Tunable Parameters
In the main paper, we have described that our model only trained for 10 epochs on the GSV-Cities [1] dataset using 3.5 hours. This is less than the full-day training time used by CosPlace/EigenPlaces, and less than the 30 training epochs used by MixVPR (also trained on GSV-Cities). To further investigate the data efficiency of our method, we reduce the training epoch (and training data), and the yielded results are shown in Table 7. When the model is trained with only 10% of the training data for 1 epoch (i.e., 0.1 epoch), our method achieves better performance than previous methods (except EigenPlaces), e.g., CosPlace and MixVPR, on Pitts30k and Tokyo24/7. The training time used is only 0.038h (i.e., 2.3min). The advantages of our method in data efficiency are mainly due to the fact that the adapter-based method maintains the powerful representation ability of the pre-trained foundation model, while our proposed cross-image encoder is an easy-to-train module. Our model has little risk of over-fitting even when trained with small amounts of data.
@@ 7 Training Epoch, Training Time, and Tunable Parameters
Table 7: The results of our CricaVPR with different training epochs.
@@
Table 9: Comparison to SOTA methods on challenging datasets. The best is highlighted in bold and the second is underlined. As in the main paper when comparing with SOTA methods, we employ PCA to reduce the descriptor dimension of our method to 4096-dim.
8 Visualizations of Place Features using t-SNE
9 Additional Results on Challenging Datasets
@@ 10 Additional Ablations on Cross-image Encoder
NPNP = 16 $\times$ 89.5 95.3 96.8 75.6 89.2 91.4 80.9 90.4 92.7
NPNP = 32 $\times$ 89.9 95.2 96.7 81.3 91.1 93.0 83.0 93.1 93.8
NPNP = 64 $\times$ 90.7 95.9 97.5 84.4 94.3 96.5 84.1 92.3 94.2
NPNP = 72 $\times$ 90.6 95.9 97.2 85.1 93.3 95.6 85.5 93.2 94.3
NPNP = 16 94.6 97.0 97.7 87.9 94.9 96.2 84.1 92.6 94.2
NPNP = 32 94.8 97.4 98.0 91.1 94.9 96.8 85.0 93.1 95.1
NPNP = 64 94.9 97.5 98.1 92.4 95.6 97.1 88.1 95.0 95.1
NPNP = 72 94.8 97.4 98.1 93.0 97.1 97.8 89.9 95.4 96.2
Table 11: Results of different training batch sizes, i.e., different numbers of places (4 images per place). NPNP is the abbreviation of “Number of Places”. We provide the results with or without the cross-image encoder.
@@
11 Effects of Training Batch Size
@@ 13 Effects of Adaptation on the Used SPM Feature
In our method, we mainly use the spatial pyramid model (SPM) representation that combines class tokens and GeM. An interesting phenomenon is that when using the frozen DINOv2 as the backbone, the SPM feature (FrozenDINOv2-SPM) performs worse than GeM (FrozenDINOv2-GeM) on Pitts30k and Tokyo24/7 (see Table 13). However, after using our adaptation, Adapt-SPM performs much better than Adapt-GeM. This shows that our adaptation makes the combined class token and GeM features in the SPM representation more compatible.

### v2

@@ 4.2 Implementation Details
We fine-tune our model on two NVIDIA GeForce RTX 3090 GPUs using PyTorch. The resolution of the input image is 224×224 and the token dimension of the backbone (ViT-B/14) is 768. Our model outputs the 14×768-dim original global features, and we use PCA for dimensionality reduction. The bottleneck ratio of our adapters is set to 0.5, so the input dimension of the three convolutional paths is 384. The 1×1 convolution before the 3×3 and 5×5 convolution reduces the channels to 24. The output dimensions of the three convolutional paths are 192, 96, and 96. The scaling factor $s$ in Eq. 4 is set to 0.2. We set the hyperparameters $\alpha=1,\beta=50,\lambda=0$ in Eq. 5 and margin = 0.1 in online mining, as in GSV-Cities [1]. We fine-tune the model using the Adam optimizer with the initial learning rate set as 0.0001 and multiplied by 0.5 after every 3 epochs. A training batch contains 72 places with 4 images each (i.e. 288 images). Training is implemented until the R@5 on Pitts30k does not improve for 3 epochs. An inference batch contains 8 images for Pitts30k and 16 images for others.
4.3 Comparison with State-of-the-Art Methods
In this section, we compare our CricaVPR with several SOTA VPR methods, mainly including six global-retrieval-based methods: NetVLAD [5], SFRS [23], CosPlace [8], GCL [38], MixVPR [2] and EigenPlaces [10]. Note that our work uses the same training dataset as MixVPR, i.e., GSV-Cities. Meanwhile, CosPlace and EigenPlaces are trained on individually constructed extra large-scale datasets, i.e., SF-XL. Both MixVPR and EigenPlaces are the latest works and represent the SOTA performance of the VPR methods based on global feature retrieval. Additionally, we also compare our approach with two excellent two-stage VPR methods (Patch-NetVLAD [26] and TransVPR [62]), which require time-consuming re-ranking using local features. The details of these methods are in Suppl. Material. Table 2 shows the quantitative results on Pitts30k, Tokyo24/7, and MSLS. Our CricaVPR uses PCA to reduce the feature dimensionality to 4096-dim (in this subsection), and achieves the best R@1/R@5/R@10 on all datasets.
@@ 4.3 Comparison with State-of-the-Art Methods
To further evaluate the performance of our method in extreme scenarios, we conduct experiments on three challenging datasets: Nordland, which exhibits seasonal changes; AmsterTime, which spans a very long time period; and SVOX, which shows extreme illumination and weather variations. The results, as shown in Table 3, demonstrate the significant superiority of our method compared to other SOTA methods. Our CricaVPR outperforms all other SOTA methods with 14.5%, 15.8%, and 20.7% absolute R@1 improvements on Nordland, AmsterTime, and SVOX-Night, respectively. This further highlights that the global image representation of our method is highly robust.
@@ 4.4 Ablation Study
Table 4: Ablation on cross-image awareness. The “+Crica” represents the addition of our cross-image correlation awareness to get the final global feature. The “SPM” represents our spatial pyramid model representation, while “SPMG” is the spatial pyramid model solely based on GeM. Except for the FrozenDINOv2-SPM that directly uses an untuned backbone (as baseline), all other versions use our adaptation method for fine-tuning.
@@
Ablation on cross-image correlation awareness. The cross-image correlation awareness achieved by the cross-image encoder after the backbone is the most important module in our method. We compare the performance of the three kinds of global features with or without across-image awareness. These features are GeM, the spatial pyramid model representation solely based on GeM (SPMG), and our spatial pyramid model representation (SPM) which uses both class token and GeM. The results are shown in Table 4. After incorporating the proposed cross-image correlation awareness (Crica), all three features achieve significant performance improvements. Due to the already impressive performance of the SPM feature after our model adaptation (AdaptSPM), the improvement provided by Crica on this feature is not as pronounced as on the GeM and SPMG features. Nevertheless, AdaptSPM+Crica still resulted in 4.2%, 7.9%, and 4.4% absolute R@1 improvements over the AdaptSPM feature on Pitts30k, Tokyo24/7, and MSLS-val, respectively. Moreover, AdaptGeM+Crica achieves an impressive 17.4% absolute R@1 improvement over AdaptGeM on Tokyo24/7. With the combined effect of our Crica and model adaptation, our method achieves nearly 2× higher R@1 on Tokyo24/7 and MSLS-val compared to the direct use of frozen DINOv2 with the SPM representation (FrozenDINOv2-SPM).
@@
Figure 6: The output feature map (attention) visualizations of pre-trained DINOv2 and adapted DINOv2. The regions attended to by pre-trained DINOv2 have no relevance to place recognition. However, adapted DINOv2 focuses on discriminative areas for VPR. Buildings that remain relatively unchanged over time receive the highest attention. Vegetation that is not expected to change in the short term receives moderate attention. Non-discriminative elements such as the sky, ground, and dynamic vehicles, are ignored.
Ablation on adaptation. We first use only the GeM features alone (without cross-image awareness) to demonstrate the performance improvement achieved by our adaptation method. As shown in Table 5, MulConvAdapter-GeM using our adaptation achieves a significant improvement over FrozenDINOv2-GeM. Especially on MSLS-val, which has more dynamic interference, our adaptation achieves nearly 2× higher R@1. Fig. 6 vividly illustrates the underlying reasons. The adapted DINOv2, in contrast to the pre-trained DINOv2, exhibits a stronger ability to focus on objects related to place recognition, with more attention given to more important objects. Table 5 also shows the performance of different fine-tuning methods when using the proposed global features (i.e. the SPM feature with cross-image correlation awareness). The FullTunedDINOv2 achieves a notable improvement over FrozenDINOv2 on Pitts30k and MSLS-val. However, because our training data has no night images like those in Tokyo24/7, FullTunedDINOv2 performs worse than FrozenDINOv2 on Tokyo24/7, i.e., it suffers from catastrophic forgetting. This indicates the necessity of parameter-efficient fine-tuning (using an adapter). Besides, ConvAdapter (as in Convpass [31]) uses 3×3 convolution to introduce local inductive biases into the model. However, it brings inappropriate local priors for VPR and results in performance degradation compared to VanillaAdapter. Our method (MulConvAdapter) uses multi-scale convolution to introduce more proper local information and thus achieves the best performance.
@@
Impact of descriptor dimensionality. In this subsection, we analyze the impact of descriptor dimensionality, and the results are shown in Table 6. Our method gets the best performance when using PCA to reduce the descriptor dimension to 4096-dim, so it is the default dimensionality we recommend. Furthermore, we continue to reduce the dimensionality to observe the point at which performance starts to noticeably decline on each dataset. For Pitts30k, the 512-dim descriptor still achieves an impressive 94.5% R@1, with no significant decrease compared to the 4096-dim descriptor. However, using the 512-dim descriptor on the other two datasets results in an obvious performance drop. This is mainly due to the drastic condition changes and the perceptual aliasing issue in these datasets, requiring higher-dimensional descriptors to provide sufficient information to distinguish places. When there is a pressing need for low-dimensional descriptors, we suggest using the 1024-dim or 2048-dim descriptor for the place images with obvious condition changes (e.g., Tokyo24/7 and MSLS), the 512-dim descriptor for images like those in Pitts30k.
Training time and data efficiency. Our model only costs 3.5 hours for training, which is significantly less than the full-day time used by CosPlace/EigenPlaces. The training epochs of ours (10 epochs) are also less than MixVPR (30 epochs) using the same dataset. To further investigate the training time and data efficiency of our method, we reduce the training epoch and training data, and the yielded results are shown in Table 7. When the model is trained with only 10% of the training data for 1 epoch (i.e., 0.1 epoch), our method achieves better performance than previous methods (except EigenPlaces) on Pitts30k and Tokyo24/7. The training time used is only 0.038h (i.e., 2.3 min). The advantages of our method in data efficiency are mainly due to the fact that the adapter-based method maintains the powerful representation ability of the pre-trained foundation model, while our proposed cross-image encoder is an easy-to-train module.
@@
Table 7: The results of CricaVPR with different training epochs.
5 Conclusions
In this paper, we presented CricaVPR, a robust global representation method with cross-image correlation awareness for VPR. Our method leverages the cross-image encoder equipped with the attention to establish the correlation among multiple images within a batch, enabling the model to harvest useful information from other images while generating the feature representation of an image. This makes the produced global features condition-invariant, viewpoint-invariant, and capable of addressing perceptual aliasing. Furthermore, we proposed a multi-scale convolution-enhanced adaptation method to introduce proper local information and effectively unleash the capability of the pre-trained foundation model for VPR. Experimental results on several VPR benchmark datasets demonstrate that our CricaVPR can provide a robust global representation to address various challenges in VPR and outperforms SOTA methods by a significant margin.
Acknowledgments
Supplementary Material
6 Overview
7. Visualizations of Place Features Using t-SNE
8. Tunable Parameters
9. Additional Results on Challenging Datasets
10. Additional Ablations on Cross-image Encoder
11. Effects of Batch Size
12. Additional Ablations on MulConvAdapter
13. Effects of Adaptation on the Used SPM Feature
14. Comparison to Other Methods with the Same
Training Dataset
15. Datasets Details
16. Compared Methods Details
17. Additional Qualitative Results and Failure Cases
18. Limitations
Note that the experiments in this supplementary material are conducted as in the main paper. That is, PCA is used to reduce the descriptor dimensionality to 4096-dim when comparing our method with other methods. However, it is not used by default in ablation experiments.
7 Visualizations of Place Features using t-SNE
8 Tunable Parameters
@@ 9 Additional Results on Challenging Datasets
Table 9: Comparison to SOTA methods on challenging datasets. The best is highlighted in bold and the second is underlined. We employ PCA to reduce the descriptor dimension of our method to 4096-dim.
@@
Moreover, we also provide the results of our CricaVPR on Pitts250k (97.5% R@1) in Section 14.
@@ 10 Additional Ablations on Cross-image Encoder
$NP$ = 16 $\times$ 89.5 95.3 96.8 75.6 89.2 91.4 80.9 90.4 92.7
$NP$ = 32 $\times$ 89.9 95.2 96.7 81.3 91.1 93.0 83.0 93.1 93.8
$NP$ = 64 $\times$ 90.7 95.9 97.5 84.4 94.3 96.5 84.1 92.3 94.2
$NP$ = 72 $\times$ 90.6 95.9 97.2 85.1 93.3 95.6 85.5 93.2 94.3
$NP$ = 16 94.6 97.0 97.7 87.9 94.9 96.2 84.1 92.6 94.2
$NP$ = 32 94.8 97.4 98.0 91.1 94.9 96.8 85.0 93.1 95.1
$NP$ = 64 94.9 97.5 98.1 92.4 95.6 97.1 88.1 95.0 95.1
$NP$ = 72 94.8 97.4 98.1 93.0 97.1 97.8 89.9 95.4 96.2
Table 11: Results of different training batch sizes, i.e., different numbers of places (4 images per place). $NP$ is the abbreviation of “Number of Places”. We provide the results with or without the cross-image encoder.
@@
11 Effects of Batch Size
@@
Batch Pitts30k Tokyo24/7 MSLS-val
Size R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
1 91.6 95.7 96.9 89.5 94.6 96.2 88.5 95.1 95.7
4 93.9 97.2 97.7 87.3 93.7 94.6 88.0 95.5 96.5
8 94.8 97.4 98.1 91.7 96.2 97.5 89.1 95.1 95.9
16 93.7 97.0 98.1 93.0 97.1 97.8 89.9 95.4 96.2
32 93.0 96.9 97.9 92.7 96.2 97.5 88.9 95.5 96.2
Table 12: Results of different inference batch size.
In addition, we also conduct experiments to study the impact of different batch sizes during inference, i.e., inference batch size. The results are as shown in Table 12. Since our method learns cross-image correlation-aware representation during training, setting the batch size to 1 during testing makes our cross-image encoder ineffective, further leading to the gap between training and testing, i.e., performance in this case will be reduced. Besides, an inference batch size that is too small (e.g., 4) will lead to unstable results (even worse than when it equals 1). Although the inference batch size that achieves the best performance on different datasets does not appear to be fixed (too small or too large will reduce performance), setting it to 16 can achieve excellent results on all datasets. So we set it to 16 (except on Pitts30k/Pitts250k we set it to 8 for better results).
@@ 13 Effects of Adaptation on the Used SPM Feature
In our method, we mainly use the spatial pyramid model (SPM) representation that combines the class token and the GeM feature. An interesting phenomenon is that when using the frozen DINOv2 as the backbone, the SPM feature (FrozenDINOv2-SPM) performs worse than GeM (FrozenDINOv2-GeM) on Pitts30k and Tokyo24/7 (see Table 14). However, after using our adaptation, Adapt-SPM performs much better than Adapt-GeM. This shows that our adaptation makes the combined class token and GeM features in the SPM representation more compatible.
14 Comparison to Other Methods with the Same Training Dataset
Most methods (except MixVPR) use different training datasets than our method. The GSV-Cities dataset used in our method has been shown to achieve better results than the datasets with weak supervision (e.g., Pitts30k and MSLS) [1]. Training different methods with the same dataset can promote fair comparisons. However, completely achieving it is hard as some methods are designed based on the characteristics of a certain (type of) dataset, and training on others may make some components of them meaningless. To minimize the impact of the training dataset on results, we use the results (reported in the MixVPR paper) of NetVLAD and CosPlace (both based on ResNet50) trained on GSV-Cities for a more fair comparison. The results are shown in Table 15 and our method still significantly outperforms others. Note that in this section we have added the results on Pitts250k (larger but easier than Pitts30k). Besides, we also provide the results of training our model on the smallest/weakest Pitts30k dataset in Table 16. Our model trained on Pitts30k still gets SOTA results (better than EigenPlaces trained on SF-XL and much better than SFRS also trained on Pitts30k).
Method Training set Pitts250k MSLS-val
R@1 R@5 R@10 R@1 R@5 R@10
CosPlace‡ SF-XL 92.3 97.4 98.4 87.4 94.1 94.9
NetVLAD† GSV-Cities 90.5 96.2 97.4 82.6 89.6 92.0
CosPlace† GSV-Cities 91.5 96.9 97.9 84.5 90.1 91.8
CricaVPR GSV-Cities 97.5 99.4 99.7 90.0 95.4 96.4
Table 15: The results of methods trained on GSV-Cities. The suffix †/‡ means that the method is different from the main paper on the backbone and/or training set. Since SF-XL is built for CosPlace (or it is part of CosPlace), CosPlace‡ trained on SF-XL is better than CosPlace† trained on GSV-Cities.
Method Training set Pitts30k Pitts250k
R@1 R@5 R@10 R@1 R@5 R@10
SFRS Pitts30k 89.4 94.7 95.9 90.7 96.4 97.6
MixVPR GSV-Cities 91.5 95.5 96.3 94.1 98.2 98.9
EigenPlaces SF-XL 92.5 96.8 97.6 94.1 97.9 98.7
CricaVPR* Pitts30k 93.0 96.9 97.9 95.9 99.0 99.5
Table 16: Results of CricaVPR* trained on Pitts30k.